Suppose that market demand can be represented as p = 100 - 2Q. There are 10 identical firms producing an undifferentiated product, each with the total cost function TC = 50 + q

2

q2

. Compare the competitive outcome with the cartel outcome. What is the individual firm's incentive to cheat on the cartel?


1
Expert's answer
2021-11-14T17:37:58-0500

"P=100-2Q"

"TC=50+Q^2"

This means that each firm has "MC=2Q"

Setting "MC=P" gives the individual firm's supply to the market:

"\implies 2Q=P"

"2Q=100-2Q"

"Q=25"